# PhysicsCorrect: A Training-Free Approach for Stable Neural PDE Simulations

### v1

@@ Abstract
Neural networks have emerged as powerful surrogates for solving partial differential equations (PDEs), offering significant computational speedups over traditional methods. However, these models suffer from a critical limitation: error accumulation during long-term rollouts, where small inaccuracies compound exponentially, eventually causing complete divergence from physically valid solutions. We present PhysicsCorrect, a training-free correction framework that enforces PDE consistency at each prediction step by formulating correction as a linearized inverse problem based on PDE residuals. Our key innovation is an efficient caching strategy that precomputes the Jacobian and its pseudoinverse during an offline warm-up phase, reducing computational overhead by two orders of magnitude compared to standard correction approaches. Across three representative PDE systems – Navier-Stokes fluid dynamics, wave equations, and the chaotic Kuramoto-Sivashinsky equation – PhysicsCorrect reduces prediction errors by up to 100× while adding negligible inference time (under 5%). The framework integrates seamlessly with diverse architectures including Fourier Neural Operators, UNets, and Vision Transformers, effectively transforming unstable neural surrogates into reliable simulation tools that bridge the gap between deep learning’s computational efficiency and the physical fidelity demanded by practical scientific applications.
@@ 1 Introduction
Simulating physical systems governed by partial differential equations (PDEs) is fundamental to numerous scientific and engineering disciplines. Achieving stable long-term rollouts is especially critical for applications such as optimal control, inverse design, and computational imaging. While classical numerical methods like finite-difference (smith1985numerical, ), finite-element (reddy1993introduction, ), and spectral-element methods (patera1984spectral, ) provide accurate solutions, they often demand substantial computational resources, limiting their real-time applicability.
Neural PDE solvers have emerged as promising alternatives offering significant computational efficiency (guo2016convolutional, ; zhu2019physics, ; geneva2020modeling, ; wandel2021teaching, ; huang2024lordnet, ; zhou2024unisolver, ; alkin2024universal, ; gao2025generative, ). These approaches approximate PDE solution operators, enabling rapid inference once trained. However, they face a fundamental challenge: error accumulation during autoregressive rollouts, where small errors compound exponentially, leading to numerical instability and divergence (Sanchez-Gonzalez2020, ).
Current mitigation strategies, such as injecting random noise during training (Sanchez-Gonzalez2020, ) or multi-step training regimes, often fall short because prediction errors are structured rather than random. Newer approaches like PDE-refiner (lippe_pde-refiner_2023, ) employ generative diffusion models at each step, but introduce substantial computational overhead that may negate the speed advantages of neural solvers.
@@ 2 Background & Related Work
The Fourier Neural Operator (FNO) (li_fourier_2020, ) established a foundation by performing spectral domain convolutions to capture global dependencies with excellent generalization. The field has since evolved with specialized architectures: message-passing networks (brandstetter_message_2022, ) for complex geometries, Clifford neural networks (brandstetter_clifford_2022, ) for physical invariances, and transformer-based approaches like GNOT (hao_gnot_2023, ), Transolver (wu_transolver_2024, ), and CViT (wang_cvit_2025, ) that combine attention mechanisms with physical priors.
For time-dependent PDEs, these models are typically applied autoregressively, predicting the next state from the current one. While they excel at one-step predictions on in-distribution data, they struggle with long-term stability. Small errors accumulate and amplify over multiple time steps, eventually causing catastrophic divergence – one of the most significant barriers to deploying neural operators in real-world applications.
@@
Several approaches have been developed to address this challenge. Sanchez-Gonzalez2020 introduced adversarial training with random noise injection to build robustness against perturbations, while brandstetter_message_2022 proposed multi-step training where loss is computed over prediction sequences, allowing models to compensate for their own errors.
Recent advances leverage generative models: PDE-Refiner (lippe_pde-refiner_2023, ) adapts diffusion models to iteratively denoise predicted states, and hu_wavelet_2024 employs wavelet-based diffusion to improve spectral accuracy. While effective, these methods introduce substantial computational overhead during training or inference – often negating the efficiency advantages of neural surrogates.
@@
An alternative strategy leverages the governing equations to correct predictions without requiring model retraining. cao_residual-based_2023 developed PDE residual minimization methods for Bayesian inference using goal-oriented a-posteriori error estimation (jha_goal-oriented_2022, ), later extended to nonlinear variational problems (jha_residual-based_2024, ). While promising, these approaches often introduce significant computational overhead.
Other research has focused on enforcing specific physical constraints: jiang_enforcing_2020 implemented spectral projection layers for divergence-free conditions in fluid simulations, and duruisseaux_towards_2024 generalized this to broader linear differential constraints. These methods are computationally efficient but limited to specific physical aspects rather than ensuring complete PDE satisfaction.
@@
Our work bridges these approaches through a correction framework that enforces PDE satisfaction at each timestep via residual minimization. Unlike existing methods, our approach requires no additional training, employs efficient caching strategies to minimize computational overhead, and generalizes across different PDE types and neural architectures. By treating physical consistency as an online correction problem we achieve stable long-term rollouts while preserving the computational advantages of neural PDE solvers.
@@ 3 Methodology
$S(\mathbf{u},\frac{\partial\mathbf{u}}{\partial t},\frac{\partial^{2}\mathbf{u}}{\partial^{2}t},\frac{\partial\mathbf{u}}{\partial\mathbf{x}},\frac{\partial^{2}\mathbf{u}}{\partial^{2}\mathbf{x}})=0,$ (1)
@@ 3.1 The PhysicsCorrect Framework: Linearized PDE Residual Correction
$L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c})\approx L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})+\frac{\partial L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})}{\partial\hat{\mathbf{u}}_{t+1}}\mathbf{u}_{t+1}^{c}.$ (5)
This approximation is valid when the correction term $\mathbf{u}_{t+1}^{c}$ is sufficiently small, which is generally
@@
the case when the neural operator has been adequately trained. Setting the linearized residual to zero yields a linear system:
@@
This can be expressed in the standard form $A\mathbf{x}=\mathbf{b}$, where $A$ is the Jacobian matrix of the PDE residual with respect to $\hat{\mathbf{u}}_{t+1}$, $\mathbf{x}$ is the correction term $\mathbf{u}_{t+1}^{c}$ we seek to determine, and $\mathbf{b}$ is the negative PDE residual $-L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$. The correction term can then be obtained by solving this linear system, typically using least-squares methods to handle potential ill-conditioning or over-determined systems. Unlike training-based approaches that learn to predict corrections from residuals, our method directly inverts the linearized PDE operator, avoiding the need for additional training data and mitigating potential distribution shifts between training and inference.
@@
Our formulation assumes that the neural operator’s prediction provides a good initial approximation, allowing the linearization to be effective. This assumption generally holds for well-trained models but may be challenged in highly chaotic systems or with poor initialization. The effectiveness of our approach across different PDEs and neural architectures is demonstrated empirically in Section 4.
@@
For this caching strategy to be effective, the PDE residual must maintain linearity with respect to $\hat{\mathbf{u}}_{t+1}$. To satisfy this requirement while preserving numerical stability, we employ a semi-implicit discretization scheme that treats linear terms (e.g., diffusion) implicitly to ensure stability, while handling nonlinear terms (e.g., advection) explicitly to preserve Jacobian constancy. For example, in the 2D Navier-Stokes equations, we implement a Crank-Nicolson scheme with implicit diffusion and explicit advection terms (Appendix A). This formulation ensures that the Jacobian remains constant across all prediction steps and initial conditions, while maintaining adequate numerical stability.
As demonstrated in Figure 3, our caching strategy achieves accuracy comparable to the non-cached version while reducing computational cost by approximately 160x. The precomputation phase for a 64x64 resolution grid requires only 8.74 seconds, after which the correction adds minimal overhead to the neural operator’s inference time (0.90 vs. 0.69 seconds for 200 time steps). Importantly, even for chaotic systems where the semi-implicit discretization yields an approximate Jacobian, the correction still substantially improves rollout stability. We observe that this approach works effectively even when applying it to systems with nonlinear residual terms, as we demonstrate with the Kuramoto-Sivashinsky equation in Section 4.3.
@@ 4 Experiments
We evaluate the PhysicsCorrect framework on three representative PDE systems that vary in complexity, dimensionality, and dynamical behavior: the 2D Navier-Stokes equations (incompressible fluid flow), the 2D wave equation (second-order hyperbolic PDE), and the 1D Kuramoto-Sivashinsky equation (fourth-order nonlinear PDE exhibiting chaotic behavior).
For each system, we test our approach with three neural network architectures: Fourier Neural Operator (FNO) (li_fourier_2020, ), UNet (Wandel2020, ), and Vision Transformer (ViT) (dosovitskiy_image_2020, ). All models are trained to generalize across different initial conditions using the Adam optimizer and L1 loss. Performance is measured by the relative L2 error between predictions and high-fidelity numerical solutions.
Our experiments are designed to address three key questions: whether the correction framework improves the stability and accuracy of long-term rollouts across different neural architectures; how effective the caching strategy is in maintaining accuracy while reducing computational overhead; and how the correction performs on systems with different levels of nonlinearity and chaotic behavior. Network architecture details and training parameters are provided in Appendix B.
@@
We first evaluate our approach on the 2D incompressible Navier-Stokes equation with a Reynolds number of 1,000 and forcing term $f(x,y)=0.1\sin(2\pi(x+y))+\cos(2\pi(x+y))$ (Appendix B.1). This system exhibits complex vorticity dynamics and is widely used as a benchmark for fluid simulation.
@@
Figure 5 shows the long-term rollout performance across different neural architectures (FNO, UNet, and ViT). All baseline models exhibit error accumulation that eventually leads to complete divergence from the reference solution. In contrast, models augmented with our physics-informed corrector maintain stable and accurate predictions throughout the entire 1,000-step simulation, regardless of the underlying architecture. The PDE residual (shown in Appendix B.1) remains consistently low for all corrected models, confirming that our approach enforces physical consistency at each time step and prevents error accumulation.
@@
Figure 5: Performance comparison of our physics-informed correction approach across different PDE systems and neural architectures. Left axis (bars): Relative L2 error of baseline models (lighter colors) versus corrected models (darker colors) for Navier-Stokes (NS), wave equation, and Kuramoto-Sivashinsky (KS) equations at the final state after long rollouts (1000 time step rollout for NS and KS equations; 100 time steps for wave equation). Right axis (line): Relative computational cost of the corrected approach compared to baseline. The correction framework consistently reduces error across all PDEs and architectures with minimal computational overhead. Detailed rollout histories are provided in the Appendix.
@@
This experiment yields two important insights: first, the quality of PDE discretization directly impacts correction accuracy; and second, even with standard discretization, our method achieves substantial error reduction (approximately 85%) relative to the theoretical optimum. These findings suggest that using finer discretization schemes for residual computation could further improve correction performance in practice.
@@
We next evaluate our approach on the 2D wave equation – a second-order linear PDE that models various physical phenomena including mechanical waves, electromagnetic waves, and seismic propagation (Appendix B.2).
@@
An interesting finding emerged during our wave equation experiments: standard first-order residual prediction (predicting $\mathbf{u}_{t+1}-\mathbf{u}_{t}$) consistently failed to capture the essential physics. To resolve this we revisited the problem through the lens of second-order dynamics, training networks to predict $\delta\mathbf{u}_{t}=\mathbf{u}_{t+1}+\mathbf{u}_{t-1}-2\mathbf{u}_{t}$ instead. This formulation resonates with the oscillatory nature of wave phenomena, providing a significantly more stable formulation. This insight highlights how the representation of physical dynamics can fundamentally shape neural network performance, even before correction mechanisms are applied.

### v2

@@ Abstract
Neural networks have emerged as powerful surrogates for solving partial differential equations (PDEs), offering significant computational speedups over traditional methods. However, these models suffer from a critical limitation: error accumulation during long-term rollouts, where small inaccuracies compound exponentially, eventually causing complete divergence from physically valid solutions. We present PhysicsCorrect, a training-free correction framework that enforces PDE consistency at each prediction step by formulating correction as a linearized inverse problem based on PDE residuals. Our key innovation is an efficient caching strategy that precomputes the Jacobian and its pseudoinverse during an offline warm-up phase, reducing computational overhead by two orders of magnitude compared to standard correction approaches. Across three representative PDE systems, including Navier-Stokes fluid dynamics, wave equations, and the chaotic Kuramoto-Sivashinsky equation, PhysicsCorrect reduces prediction errors by up to 100× while adding negligible inference time (under 5%). The framework integrates seamlessly with diverse architectures, including Fourier Neural Operators, UNets, and Vision Transformers, effectively transforming unstable neural surrogates into reliable simulation tools that bridge the gap between deep learning’s computational efficiency and the physical fidelity demanded by practical scientific applications. The code is available at
https://github.com/summerwine668/PhysicsCorrect.
@@ 1 Introduction
Simulating physical systems governed by partial differential equations (PDEs) is fundamental to numerous scientific and engineering disciplines. Achieving stable long-term rollouts is especially critical for applications such as optimal control, inverse design, and computational imaging. While classical numerical methods like finite-difference (Smith 1985), finite-element (Reddy 1993), and spectral-element methods (Patera 1984) provide accurate solutions, they often demand substantial computational resources, limiting their real-time applicability.
Neural PDE solvers have emerged as promising alternatives offering significant computational efficiency (Guo, Li, and Iorio 2016; Zhu et al. 2019; Geneva and Zabaras 2020; Wandel, Weinmann, and Klein 2021; Huang et al. 2024; Zhou et al. 2024; Alkin et al. 2024; Gao, Kaltenbach, and Koumoutsakos 2025). These approaches approximate PDE solution operators, enabling rapid inference once trained. However, they face a fundamental challenge: error accumulation during autoregressive rollouts, where small errors compound exponentially, leading to numerical instability and divergence (Sanchez-Gonzalez et al. 2020).
@@
Current mitigation strategies, such as injecting random noise during training (Sanchez-Gonzalez et al. 2020) or multi-step training regimes, often fall short because prediction errors are structured rather than random. Newer approaches like PDE-refiner (Lippe et al. 2023) employ generative diffusion models at each step, but introduce substantial computational overhead that may negate the speed advantages of neural solvers.
@@ 2 Background & Related Work
The Fourier Neural Operator (FNO) (Li et al. 2020a) established a foundation by performing spectral domain convolutions to capture global dependencies with excellent generalization. The field has since evolved with specialized architectures: message-passing networks (Brandstetter, Worrall, and Welling 2022) for complex geometries, Clifford neural networks (Brandstetter et al. 2022) for physical invariances, and transformer-based approaches like GNOT (Hao et al. 2023), Transolver (Wu et al. 2024), and CViT (Wang et al. 2025) that combine attention mechanisms with physical priors.
For time-dependent PDEs, these models are typically applied autoregressively, predicting the next state from the current one. While they excel at one-step predictions on in-distribution data, they struggle with long-term rollout stability. Small errors accumulate and amplify over multiple time steps, eventually causing catastrophic divergence, one of the most significant barriers to deploying neural operators in real-world applications.
@@
Several approaches have been developed to address this challenge. Sanchez-Gonzalez et al. (Sanchez-Gonzalez et al. 2020) introduced adversarial training with random noise injection to build robustness against perturbations, while Brandstetter et al. (Brandstetter, Worrall, and Welling 2022) proposed multi-step training where loss is computed over prediction sequences, allowing models to compensate for their errors.
Recent advances leverage generative models: PDE-Refiner (Lippe et al. 2023) adapts diffusion models to iteratively denoise predicted states, and Hu et al. (Hu et al. 2024) employs wavelet-based diffusion to improve spectral accuracy. While effective, these methods introduce substantial computational overhead during training or inference, often negating the efficiency advantages of neural surrogates.
@@
An alternative strategy leverages the governing equations to correct predictions without requiring model retraining. Cao et al. (Cao et al. 2023) developed PDE residual minimization methods for Bayesian inference using goal-oriented a-posteriori error estimation (Jha and Oden 2022), later extended to nonlinear variational problems (Jha 2024). While promising, these approaches often introduce significant computational overhead.
Other research has focused on enforcing specific physical constraints. Jiang et al. (Jiang et al. 2020) implemented spectral projection layers for divergence-free conditions in fluid simulations, and Duruisseaux et al. (Duruisseaux et al. 2024) generalized this to broader linear differential constraints. These methods are computationally efficient but limited to specific physical aspects rather than ensuring complete PDE satisfaction.
@@
Our work bridges these approaches through a correction framework that enforces PDE satisfaction at each timestep via residual minimization. Unlike existing methods, our approach requires no additional training, employs efficient caching strategies to minimize computational overhead, and generalizes across different PDE types and neural architectures. By treating physical consistency as an online correction problem, we achieve stable long-term rollouts while preserving the computational advantages of neural PDE solvers.
@@ 3 Methodology
$S\left(\mathbf{u},\frac{\partial\mathbf{u}}{\partial t},\frac{\partial^{2}\mathbf{u}}{\partial^{2}t},...,\frac{\partial\mathbf{u}}{\partial\mathbf{x}},\frac{\partial^{2}\mathbf{u}}{\partial^{2}\mathbf{x}},...\right)=0,$ (1)
@@ 3.1 The PhysicsCorrect Framework: Linearized PDE Residual Correction
$L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1}+\mathbf{u}_{t+1}^{c})$ $\approx L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$
$+\frac{\partial L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})}{\partial\hat{\mathbf{u}}_{t+1}}\mathbf{u}_{t+1}^{c}.$ (5)
This approximation is valid when the correction term $\mathbf{u}_{t+1}^{c}$ is sufficiently small, a condition typically
@@
satisfied when the neural operator has been adequately trained for direct full physical prediction (directly approximate $\mathbf{u}_{t+1}$). When using the residual prediction ($\mathbf{u}_{t+1}-\mathbf{u}_{t}$), this assumption will be relaxed (see the sensitivity analysis in the Appendix). Setting the linearized residual to zero yields a linear system:
@@
This can be expressed in the standard form $A\mathbf{x}=\mathbf{b}$, where $A$ is the Jacobian matrix of the PDE residual with respect to $\hat{\mathbf{u}}_{t+1}$, $\mathbf{x}$ is the correction term $\mathbf{u}_{t+1}^{c}$ we seek to determine, and $\mathbf{b}$ is the negative PDE residual $L_{\text{PDE}}(\mathbf{u}_{t},\hat{\mathbf{u}}_{t+1})$. The correction term can then be obtained by solving this linear system, typically using least-squares methods to handle potential ill-conditioning or over-determined systems. Unlike training-based approaches that learn to predict corrections from residuals, our method directly inverts the linearized PDE operator, avoiding the need for additional training data and mitigating potential distribution shifts between training and inference.
@@
Our formulation assumes that the neural operator’s prediction provides a good initial approximation, allowing the linearization to be effective. This assumption generally holds but may be challenged in highly chaotic systems or with poor initialization (see sensitivity analysis in the Appendix). Nevertheless, the effectiveness of our approach across different PDEs and neural architectures is demonstrated empirically in Section 4, even in challenging regimes.
@@
For this caching strategy to be effective, in the Jacobian calculation, the PDE residual must maintain linearity with respect to $\hat{\mathbf{u}}_{t+1}$. To satisfy this requirement while preserving numerical stability, we employ a semi-implicit discretization scheme that treats linear terms (e.g., diffusion) implicitly to ensure stability, while handling nonlinear terms (e.g., advection) explicitly to preserve Jacobian constancy. For example, in the 2D Navier-Stokes equations, we implement a Crank-Nicolson scheme with implicit diffusion and explicit advection terms (more details are in the Appendix). This formulation ensures that the Jacobian remains constant across all predictions and initial conditions, while maintaining adequate numerical stability.
As demonstrated in Figure 3, our caching strategy achieves accuracy comparable to the non-cached version while reducing computational cost by approximately 160x. The precomputation phase for a 64x64 resolution grid requires only 8.74 seconds, after which the correction adds minimal overhead to the neural operator’s inference time (0.90 vs. 0.69 seconds for 200 time steps). Importantly, even for chaotic systems where the approximation to the Jacobian matrix is relatively poor with the semi-implicit discretization, the correction still substantially improves rollout stability. We observe that this approach works effectively even when applying it to chaotic systems with nonlinear residual terms, as we demonstrate with the Kuramoto-Sivashinsky equation in Section 4.3.
Figure 4: Performance comparison of our physics-informed correction approach across different PDE systems and neural architectures. Left axis (bars): Relative L2 error of baseline models (lighter colors) versus corrected models (darker colors) for Navier-Stokes (NS), wave equation, and Kuramoto-Sivashinsky (KS) equations at the final state after long rollouts (1000 time step rollout for NS and KS equations; 100 time steps for wave equation). Error bars represent the standard deviation for 5 seeds. Right axis (line): Relative computational cost of the corrected approach compared to baseline. The correction framework consistently reduces error across all PDEs and architectures with minimal computational overhead. Detailed rollout histories and comparison with baseline models are provided in the Appendix.
@@ 4 Experiments
We evaluate the PhysicsCorrect framework on three representative PDE systems (Figure 4) that vary in complexity, dimensionality, and dynamical behavior: the 2D Navier-Stokes equations (incompressible fluid flow), the 2D wave equation (second-order hyperbolic PDE), and the 1D Kuramoto-Sivashinsky equation (fourth-order nonlinear PDE exhibiting chaotic behavior). For each system, we test our approach with three neural network architectures: Fourier Neural Operator (FNO) (Li et al. 2020a), UNet (Wandel, Weinmann, and Klein 2020), and Vision Transformer (ViT) (Dosovitskiy et al. 2020). All models are trained to generalize across different initial conditions using the Adam optimizer and L1 loss. Performance is measured by the relative L2 error between predictions and high-fidelity numerical solutions.
Our experiments are designed to address three key questions: whether the correction framework improves the stability and accuracy of long-term rollouts across different neural architectures; how effective the caching strategy is in maintaining accuracy while reducing computational overhead; and how the correction performs on systems with different levels of nonlinearity and chaotic behavior. Network architecture details and training parameters are provided in the Appendix.
@@
We first evaluate our approach on the 2D incompressible Navier-Stokes equation with a Reynolds number of 1,000 and forcing term $f(x,y)=0.1\sin(2\pi(x+y))+\cos(2\pi(x+y))$ (more details are in the Appendix) This system exhibits complex vorticity dynamics and is widely used as a benchmark for fluid simulation.
@@
Figure 4 shows the long-term rollout performance across different neural architectures (FNO, UNet, and ViT). All baseline models exhibit error accumulation that eventually leads to complete divergence from the reference solution. In contrast, models augmented with our physics-informed corrector maintain stable and accurate predictions throughout the entire 1,000-step simulation, regardless of the underlying architecture. The PDE residual (more details and comparison with baseline models are shown in the Appendix) remains consistently low for all corrected models, confirming that our approach enforces physical consistency at each time step and prevents error accumulation.
@@
This experiment yields two important insights: first, the quality of PDE discretization directly impacts correction accuracy; and second, even with standard discretization, our method achieves substantial error reduction (approximately 85%) relative to the theoretical optimum. These findings suggest that using finer discretization schemes for residual computation could further improve correction performance in practice (further discussion is in the Appendix).
@@
We next evaluate our approach on the 2D wave equation, a second-order linear PDE that models various physical phenomena, including mechanical waves, electromagnetic waves, and seismic propagation (more details are in the Appendix).
@@
An interesting finding emerged during our wave equation experiments: standard first-order residual prediction (predicting $\mathbf{u}_{t+1}-\mathbf{u}_{t}$) consistently failed to capture the essential physics (see Appendix for more results). To resolve this, we revisited the problem through the lens of second-order dynamics, training networks to predict $\delta\mathbf{u}_{t}=\mathbf{u}_{t+1}+\mathbf{u}_{t-1}-2\mathbf{u}_{t}$ instead. This formulation resonates with the oscillatory nature of wave phenomena, providing a significantly more stable formulation. This insight highlights how the representation of physical dynamics can fundamentally shape neural network performance, even before correction mechanisms are applied.